The operating potential difference of a light bulb is 120 V. The power rating of the bulb is 90 W.

(a) Find the current in the bulb. A

(b) Find the bulb's resistance.
Answer #1

a)

P = Vi

90 = 120 i

i = 0.75 A

========

b).

R = V/i = 120 / 0.75

R = 160 ohm

========

Comment in case any doubt, will reply for sure . Goodluck